Project S has a cost of $10,000 and is expected to produce benefits (cash flow) of $3,000 per year for 5 years. Project L costs $25,000 and is expected to produce cash flow of $7,400 per year for 5 years. Calculate the two projects’ NPV, IRR, and MIRR, assuming a cost of capital of 12%. If these are two mutually exclusive projects, which project would be selected? Justify your answer(s).   

HAND WRITE WORK PLEASE :)
Answer #1

NPV of Project S =PMT*((1-(1+r)^-n)/r)-Cost=3000*((1-(1+12%)^-5)/12%)-10000 =814.13
NPV of Project L =PMT*((1-(1+r)^-n)/r) -Cost=7400*((1-(1+12%)^-5)/12%)-25000 =1675.34

IRR of Project S using Financial Calculator
N=5;PMT =3000;PV=-10000;CPT I/Y =15.24%

IRR of Project L using Financial Calculator
N=5;PMT =7400;PV=-25000;CPT I/Y =14.67%

FV of Cash Flows of S =PMT*((1+r)^n-1)/r) =3000*(((1+12%)^5-1)/12%)=19058.5421
MIRR =(FV/PV)^(1/n)-1 =(19058.5421/10000)^(1/5)-1 =13.77%

FV of Cash Flows of L =PMT*((1+r)^n-1)/r) =7400*(((1+12%)^5-1)/12%)=47011.07
MIRR =(FV/PV)^(1/n)-1 =(47011.07/25000)^(1/5)-1 =13.46%
​​​​​​​
Project L should be selected as it has higher NPV


​​​​​​​